Act ^{of} War

Version 1.1 A Game by Dave Smith ^{∗} csc361@cent1.lancs.ac.uk

September 1992

∗Manual LaTEX’ed by Thomas Bätzler, BAETZLER@IRAV17.IRA.UKA.DE

Act of War User’s Manual 2

Contents 1 System requirements

Act Of War works under both Workbench 1.3 and 2.0, in both NTSC and PAL, and requires 1 meg of memory as a minimum. The game may need more than a meg to run it from a hard drive.

It should work fine on accelerated systems, if a little fast in some places! I’ve got some “slowdown” routines in it, but as I’ve only got a 68000 I haven’t been able to test it.

As Act Of War was written with AMOS, it multitasks - use Left-Amiga A to switch between the game and Workbench.

2 Introduction

Act of War, or AoW as I’ve just decided to call it as it involves less typing, is a strategy game in the tradition of Laser Squad, Breach 2 (or so I’m told) and, if anyone can remember that far back, Rebelstar.

If you’ve never played any of those games, think of a childish fascination for big guns married to a fair amount of strategy and you won’t be far wrong.

2.1 So what makes AoW different?

The game is played over a series of missions, three of which are included with the program. More missions are being worked on as you read this, by myself and others. And if you don’t like these missions, there’s no reason for you not to create your own!

On payment of a small shareware fee (more information later on) you’ll get the mission designer that I use, which makes mission creation very simple. Everything from the map, to the weapons, to the sound effects can be changed.

2.2 Noteable game features include:

• 1 or 2 players

• Sampled sound

• Completely extendable!

• 3 difficulty levels

• Optional line-of-sight feature

• Squad arming sequence

• Explosive scenery • Keyboard and / or mouse controlled

• Halfway-decent user interface

• Revolutionary new ICASS AI system (see later)

The three missions included in this release are as follows:

Informant — An ex-director of your shady corporation is ready to spill your industrial secrets to a rival company. There’s only one way to stop him - break into his sprawling mansion and take him out! (Obviously, negotiation comes a long way down the list with these people).

Escape — Taken prisoner by a corrupt Planetary Security Force for your part in the above escapade, you and two companions have been jailed in an orbital maximumsecurity prison. Initially armed with only a little smuggled-in plastic explosive, you must break out of your cell, somehow obtain a weapon, free your comrades and escape to a passing PSF shuttle to secure your freedom.

Terminator! — Soon after you make it back to your base after escaping captivity, a Terminator droid from a rival corporation stages an assault! Backed up by a squad of combat droids, it will attempt to kill you all …

3 Installation

AoW can be installed on a hard drive - just make sure that the files:

diskfont.library mathtrans.library

are in the libs: drawer, and that you’ve copied the “Spacey” font to the fonts: drawer. The directory that you put the game in must have the name “px:” assigned to it. In other words, if the game is in “dh1:Games/AoW”, you should put:

assign px: dh1:Games/AoW

in your startup-sequence.

The mission directories should be in the same dir as the game, ie. “px:”. The default missions also require that the “Pics” drawer and the “Snd” drawer, and their contents, are present in “px:” also.

The game is much nicer installed on a hard drive (isn’t everything?). As it is so extendable, a fair amount of disk access is inevitable. However, there is no disk access during a game so the actual gameplay speed is identical on both floppy and hard drive systems.

To summarise then: the necessary files are as follows. libs:diskfont.library

mathtrans.library fonts:Spacey/7

Spacey.font px:Pics/GameTiles EscapeTiles GameBack.iff WeaponBack.iff Weapons.1

Snd/Samples.1 Act Of War Informant/Everything Escape/Everything Terminator!/Everything

4 Gameplay

I know, I know, you just want to know how to play!

4.1 The front end

• The first thing you have to do is load a mission in. Select “Load Mission” from the menu. The disk will grind a bit, and a list of the available missions will come up. Choose the one you want, again using the mouse.

• Select the difficulty level you want. Easy gives you a lot of cash for the arming sequence (see below). Medium gives you less cash and introduces the line-of-sight feature, where you can only see enemy units that are in your line of sight. Hard keeps line-of-sight and gives you the least cash.

• Choose a 1 or 2 player game. The AI is not brilliant, but I think it’s good enough to be called “not bad”. It even has its own acronym, which is obviously essential for any program these days. So, when you play a 1-player game of AoW, you are experiencing ICASS - the Intellectually Challenged Artificial Stupidity System (tm).

Now, don’t you feel proud?

4.2 Arming

• Next comes the arming sequence. A total amount of cash is shown, and you can either accept the default weapons (use the “Next Unit” icon, which looks like a tick, to see each unit’s weapons) or choose to buy your own. To keep the defaults, click the “Quit” icon; to buy weapons, click either of the two “Buy” buttons.

• The squad’s default weapons are usually round about what you could buy at the

Medium level of difficulty. The amount of cash you get at each level is as follows:

Easy  — 250 per unit
Medium — 200 per unit
Hard  — 175 per unit

• If you do buy new weapons, the first button is “Buy Weapon”, the next is “Buy Ammo”, the two arrow icons let you see the weapons and the “Next Unit” button is used when you’ve finished arming that unit. You can always go back to it, until you quit the arming.

• The cash shown is a total for your whole squad, try to divide weapons fairly equally. It’s always better to have, say, four men with Pulse rifles rather than one man with a Plasma Cannon and tons of ammo.

• If you buy something, then decide you don’t want it, you can sell it back at no loss by clicking on the object’s name in the inventory.

• Click on the Quit button (bottom right) when you’ve finished.

4.3 The game

• The screen is divided into four windows:

1. The largest window, at the top left, is the view window. This is where the action happens.

2. Directly below this is the message window. Always keep half an eye on this for reports and updates.

3. At the top left is the status window. Here you can see the stats of the current unit.

4. The final window at the bottom right is the mini-map window. Apart from giving you an idea of the overall shape of the map, you can keep track of your units here via the flashing dots which show their positions.

• Player 1 always moves first. You begin the game fully armed and ready.

• There are several ways of moving around the map. You can use the cursor keys or the joystick to move the cursor around. You can click with the left mouse button on the view window to place the cursor directly onto a square. You can click on the mini-map to jump directly to another area.

• To select a character, press the joystick fire button or the space key when the cursor is over a unit. You can also click the right button on a unit to select it. You are now in move mode, where believe it or not you get to move the unit around.

• To move a unit, use either the cursor keys, the keypad or the joystick. The keypad allows movement in 8 directions, using the keys:

7 8 9 4 6 1 2 3

• To deselect a character, use the space key, the joystick fire button or the right mouse button at any time.

• If you are playing on Easy level, you’ll be able to see all the enemy units as you look around the map. This takes a lot of the fun out of the game, so the default difficulty level is Medium. Here, you can only see an enemy if they are in a direct line-of-sight. This allows them to hide around corners and jump out at you, wait behind doors etc. Remember, you can see through windows.

• If you find a scenario too tough on Medium, or you don’t like the hidden movement, go to Easy. Otherwise, I’d recommend you stay on Medium or Hard.

So now you know how to move around the map, and select characters (by the way - I use the words “character” and “unit” interchangeably, they both mean the same). Now things get a little bit more complicated!

Nearly every action in the game can be carried out by the mouse or the keyboard. If you prefer to use a keyboard, you can play the whole game without having to take your hands away from it. However, I prefer a mix of the mouse and keyboard. It’s up to you though.

You’ll have noticed (assuming you’ve actually played the game yet) six buttons by the mini-map window. I’ll go through these and any sub-menus that they lead to.

First, a word about actions. Everything you do in the game requires a certain amount of actions to carry out, and you have a set number of actions per unit, per turn. For instance, moving across a normal square uses 2 actions, but moving over rubble and debris takes 3. Picking something up takes 5 actions, teleporting takes 20. As I explain each option, the number of actions it needs is shown alongside.

If a character’s health is reduced to 1/3 of his (or her!) maximum, they are said to be injured and their actions are reduced by half. You can tell when a character is injured - their health and actions show up red in the status window.

MOVE  Keyboard: SPACE^{1}  Action needed: —

The first button, with the four outwardly-pointing arrows, is the MOVE button. When a unit is in FIRE mode, this button exits and returns it to the default MOVE mode.

FIRE  Keyboard: F  Action needed: —

The gunsight icon begins FIRE mode if a unit is selected. This changes the cursor to a gunsight, and brings up a sub-menu in the mini-map window (all the sub-menus appear here).

The cursor can be moved in the same way that characters are moved around. When you’ve moved it over a target, select a type of shot from the menu, using the mouse, or one of the keyboard short-cuts:

Aimed shot  Keyboard: A  Action needed: varies

Fires an aimed shot towards the gunsight. The number of action points needed is shown.

Fast shot  Keyboard: F  Action needed: varies

Fires a fast shot towards the gunsight. Actions needed vary according to the weapon you’re using. A Fast shot requires only 2/3 of the action points needed for an Aimed shot, but has only 2/3 the chance of being accurate. This doesn’t sound very inaccurate, but in practice you’ll find it’s more than enough to miss rather a lot!

Throw Grenade  Keyboard: G  Action needed: 6

If a character has a grenade, throw one toward the cursor. The cost is always 6 actions. Grenades have several special properties. Naturally, they are explosive, but where a shot will be stopped by a window, a grenade will continue through it. Where a shot will hit a desk or other piece of furniture, a grenade can be chucked over it. Where a shot will hit a character between you and your target, a grenade can be thrown past them. Grenades will always head towards where you threw them - there is no such thing as an off-target grenade. There is a downside to this. Grenades are not the Ultimate Weapon! A grenade can only be thrown a finite distance, unlike the other weapons which have no maximum range. A very strong character, such as a Battle Droid, can throw one maybe 15 squares (think of this as about 30 metres). An average human can manage about 10 squares. They are also expensive for a one shot weapon.

Load Weapon  Keyboard: L  Action needed: varies

Assuming you have any ammo, this loads your current weapon. Actions needed vary according to the weapon you have.

Weapons are divided into 2 basic categories: explosive and non-explosive. Explosive weapons kill outright, unless the target is incredibly tough. They also affect the environment around the target, eg. walls may be knocked down, windows smashed, other units injured, volatile substances ignited etc. They are very powerful, and you’ll probably find that they’re priced accordingly!

Non-explosive weapons, the majority, simply damage characters. Armour worn by a character affects how much damage they sustain. Beware of firing any type of weapon near things like oil drums, petrol tanks, ammo crates etc. Even non-explosive weapons set them off! Windows, too, can be smashed by any type of weapon.

Hand-to-hand weapons are also available, though the only one available in the three included scenarios is the Laser Knife. These are used purely for hand-to-hand combat.

This type of combat is initiated during a character’s movement, by walking into an enemy unit. The attacker always has the advantage, but if the enemy is strong and skilful, it is possible for the aggressor to be injured rather than his target. You always have to be careful during hand-to-hand combat! If possible, make sure your enemy is less suited to it than you.

CENTRE VIEW  Keyboard: C  Action needed: —

This button centres the view screen on the square currently occupied by the cursor. You can recognise it by its four inwardly-pointing arrows.

OPTIONS MENU Brings up a sub-menu in the mini-map screen.

Which menu you get depends on whether you have a character currently selected or not. If you don’t, you get the standard menu:

Next Unit  Keyboard: N  Action needed: —

Moves the view to the next selectable unit.

Success Rate  Keyboard: P  Action needed: —

Each game is won on the destruction of all enemy units, or the accumulation of success points. You are always told when your success rating goes up, and in the documentation for each individual mission you should be told exactly what gets your success up. It can be reaching a particular place, or blowing it up, or killing a particular enemy. The game is won when you or your opponent has a 100 Anyway! This option shows your current rate in the message window.

Turns Left  Keyboard: T  Action needed: —

Some missions have a time limit. Selecting this option shows you how long you’ve got left, or tells you “Unlimited time available”.

Save Game  Keyboard: Shift-S  Action needed: —

This brings up a file selector. Choose the name you want to save the game as, and press RETURN or choose OK with the mouse. Each save game is approximately 10-12k long. A few notes on the AMOS file requestor: at first glance, it appears to have no buttons for “Parent” or to bring up a list of active drives and assigns. Actually, it can do both things, in its own cute but slightly eccentric way.

To get to the parent of a directory, click on the tiny circle icon in the top-leftish corner, above the two microscopic arrows. To get a list of active drives, use the right mouse button.

Load Game  Keyboard: Shift-L  Action needed: —

Loads a game, via the same requestor as above.

Quit Game  Keyboard: Shift-Q  Action needed: —

Quits a game. There’s no “Sure?” type warning, so make sure you really want to quit before you choose this!

The other menu is called up when you have a character selected. It is from this menu that most of the interesting actions in the game can be accessed from.

Inventory  Keyboard: I  Action needed: —

Brings up a display of the items in your inventory. Space or selecting CANCEL quits (note: you can’t select CANCEL if you called the inventory up via the keyboard. If you start an option with the keyboard, you can’t use the mouse with it. This applies to all the options).

Success Rate  Keyboard: P  Action needed: —

Shows your current success rate in the message window.

Pick Up  Keyboard: G  Action needed: 4

Pick up something from the ground. When you’ve shot and killed an enemy, you can, if you like, stand over his body and unashamedly loot it. This is often an essential tactic! All’s fair in love and war (although it’s often considered rude if you kill someone and loot the corpse if love is concerned). Objects can also be lying around - they are usually well marked on the map by a chest, locker, load of objects scattered around etc. Just seeing what is in a square costs nothing. It’s only if you actually get something that your actions are lowered.

Drop  Keyboard: D  Action needed: 4

Lets you drop an object from your inventory. Note: any weapon you drop will have its ammo reduced to 0. This is NOT a bug, it is a “feature”. Although I admit, it’s a pretty bloody stupid feature.

Load Weapon  Keyboard: L  Action needed: varies

Loads your weapon, if you have any spare ammo.

Change Weapon  Keyboard: W  Action needed: 5

If you have a weapon, this swaps your current weapon for one in your inventory (you choose it, of course). If you have no weapon, you choose one to use. Note: grenades and plastic explosives don’t need to be held as weapons to be used, they just need to be in your inventory.

Next Flips to the next option screen.

Open Door  Keyboard: O  Action needed: 2

Opens a door adjacent to you. If there is more than one, you can choose the correct one by clicking on the arrow pointing to the door. If you have to click on an arrow, you can use the keyboard instead: the keypad keys work like this:

7 8 9 4 6 1 2 3

If you have an A600 without the keypad, you can use the normal number keys with the same effect. This technique works with all the menus with arrows to click on, such as Shut Door and Use Explosive (see later). If you’re going to walk through a door, there’s no need to open it first. Just walk into it, it’ll open for you.

Shut Door  Keyboard: S  Action needed: 2

Operates identically to “Open Door”, except in one vital respect (it closes them).

Teleport  Keyboard: T  Action needed: 20

If you stand on a teleport, this option operates it. It uses a large amount of actions, as it’s a bit draining to be split up into your component molecules and stuck back together several hundred metres away. Or so I’ve been told.

Use Medikit  Keyboard: M  Action needed: 8

Medikits restore a unit to full health. If a character is injured, the injury is healed also. Of course, you need to possess a medikit to be able to use one.

Prime Explosive  Keyboard: X  Action needed: 6

If you have any plastic explosive, this option allows you to set the timer on it. 1 turn means detonation at the end of this turn, 2 means at the end of the enemies’ turn, 3 means at the end of your next turn, 4 means at the end of the enemies’ next turn but one etc. Click on Increase or use the “+” key to increase, Decrease or the “–” key to reduce the time. The SPACE key or selecting OK finishes. Make sure you have enough actions to get rid of the explosive once you’ve set it!

Use Explosive  Keyboard: U  Action needed: 6

Plants one bit of explosive you have primed. Click on an arrow to place it. Placing more explosive in one area does not increase the size of the explosion.

Previous Click on this to go back to the first option screen.

REDRAW MAP

After you’ve finished with the Options menu, you may want to see the map again. This button carries this out.

END TURN

The final button ends your turn, allowing the computer or Player 2 to have their go.

That’s the basic gameplay. There’s room for a decent amount of strategy in there, and of course lots of explosions. I mean, let’s get our priorities right.

5 ShareWare

As it says when you load the game, Act Of War is a SHAREWARE product, though I bet you don’t need it shouted at you.

AoW has taken me nearly four months of practically full-time effort to write. Assuming a professional programmer is paid £10K a year, I make that £3 grand you lot owe me!

Don’t worry, just joking.

However, if you like the game, I would like a minimum of £5, which isn’t too much really (better than £3000 anyway). If you really like the game, I’d very much appreciate it if you sent more. If you think it’s commercial quality, you could send £10 or £15, or even more if you’ve got more money than sense. Of course, you get benefits yourself!

A “donation” (polite word for payment) of £5 will get you a nice letter and a disk containing the mission editor, the latest version of the program and any extra missions I’ve got. £10 will get you a very nice letter, the mission editor disk, and the source code of the program (all 275K of it) on a second disk. If you send me more than £10, you’ll get an unspeakably nice letter, all the above, and probably Christmas cards for life as well. In addition, I’ll nominate you for “Best Supporting Actor” at next year’s Academy Awards, and make sure your house is looked after when you go abroad. If you know what I mean.

I’d like the money in either:

• A British cheque or postal order, or

• UK Pounds Sterling (cash) at your own risk.

When I tried to cash some US cheques recently, I was told the minimum charge would be £10. So, it has to be UK currency I’m afraid, unless you make me an offer I can’t refuse (like a $500 cheque, or your DAT player, or your sister).

Now all you need to know is where to send the money!

From now (September 1992) until June 1993, with gaps for Christmas and Easter, you can reach me at the following address:

98 Thornton Road, Morecambe, England LA4 5PJ

During the gaps (approximately December 15th - January 15th, and March 15th - April 15th) and after June 1993, you can reach me at:

4 Cleveland View, South Bents, Sunderland England SR6 8AP

6 Updates

This is version 1.0 of the program (if it was a Mac program it would be version 1.0.0). I’ve tested it as much as possible, to the extent that I’m sick to death of it, so it should be as bug free as possible. If you find any bugs, or have any suggestions for a future version, please let me know when you register.

There’s no reason why I can’t do a V2.0 of the game, so any suggestions will be gratefully received.

The most fun part of the game is of course mission designing, so I’m working on a few more. I’ll distribute these in a couple of ways:

• Upload them to FTP sites.

If you don’t know what an FTP site is, you probably can’t reach one! Basically, an FTP site is an on-line software library, usually allowing anonymous access. They can be reached over the Internet/JANET, so if you’re at University you can probably use the mainframe to use them.

My preferred sites are:

amiga.physik.unizh.ch (130.60.80.80)

wuarchive.wustl.edu (I forget the number)

• When I’ve got another diskful, I’ll send it to UK PD libraries.

Hopefully, soon other people will be doing missions too, so if the game takes off you should be seeing a lot of missions eventually!

Missions that I and others are working on at the minute:

• Firefight - A fight in an ammo factory

• Arena - Combat in an alien arena

• Meltdown - Destroy a reactor, and escape before it blows

• Supergun - Break into a complex and steal a new experimental weapon, then use it to get out

• Zoo - Alien exhibits have escaped!

• Haunted - Destroy a mansion populated by werewolves, vampires, ghouls... lots of silver bullets and stakes here!

• Disney - You know as much as I do!

• Fantasy - Wizards, elves, orcs etc.

There’s only one way to make sure the game takes off... register!

As for my next project... does anyone remember a game called “Krakatoa” on the Spectrum? This was released in about 1983 or 84, it was reviewed in the first issue of Crash magazine anyway. It was stunning! You were this helicopter pilot, who had to protect a village and an oil tanker from a huge volcano. Also, there was an enemy who launched rocket attacks every so on.

You could do all these brilliant things like bomb the oil tanker until some survivors jumped into a lifeboat, rescue them from the lifeboat, then fly them to the volcano and drop them in! This set off the volcano, which erupted all over this innocent little village, sending people running from their homes, ready to be rescued, or bombed, or landed on, or shot, or drowned, or dropped from a great height, or lowered into a volcano.

It really was something else. Anyway, I’m planning to do a version of this. I don’t know who the original authors were, but if you’re by any chance reading this, let me know. And if you want to see a game like that on the Amiga, register for AoW so I have an incentive!

7 Acknowledgements

It’s impossible to do something like this entirely on your own, and so I’ve got a few people to thank.

First, my girlfriend Louise for putting up with endless boring monologues on the difficulties of AI programming and the benefits of so many enormous guns when faced with a rampant Battle Droid.

Gary Dietachmayer helped an enormous amount with the AI, without his brilliant ideas this may well have ended up two-player only. Oh, and he was incredibly patient with my useless terminal and its failed attempts at uploading files (not to mention the frequent mis-spelling of his name).

Martyn Brown, of Team 17 Software, granted permission for me to use several samples from the brilliant game “Alien Breed”. Listen out for that “Load Weapon” sound. Best sample I ever heard!

Nigel Mansell for winning the British Grand Prix (never mind the World Championship), giving me a fantastic day out even if I did have to stand from 4am until 4pm to see it.

This game was brought to you with the aid of:

•

Bruce Springsteen — Just about everything
REM  — Out Of Time, Green
Cyndi Lauper  — She’s so Unusual, Night to Remember
Gerry Rafferty  — Baker Street (what else?)
The Eagles  — Best Of

• Enough Coke to float a cruise liner (that’s the drink, by the way)

• AMOS and Compiler, both versions 1.34, from Europress Software DPaint III

• Boredom staved off with Civilization, which also gave me the inspiration for the “Centre view” function and the keyboard shortcuts.

• And of course, inspiration by the fantastic Rebelstar / Laser Squad.

8 Various rubbish

If you’re the sort of person who likes useless facts, you might be interested in this bit.

• The source code for this program is 275K long.

• Approximately 55K of this is sound and graphics data, the rest is just text.

• It needs a 220K text buffer in the AMOS editor to load it all in.

• The variable buffer is 150K! (The default is 4K)

• AMOS takes 34 seconds, after pressing the “Run” button, to actually run the program. The delay occurs while it tests the code.

• Saved as ASCII, it needs 95 sides of A4 to print it out.

• The hardware setup I used consists of:

– Amiga A500+ with 2MB – ICD IDE hard drive + interface

– External floppy

• The program was begun on April 15th, and finished on September 1st, 1992.

9 Background information

As I mentioned at the start of this documentation, AoW is “in the tradition of” Rebelstar and Laser Squad. I understand that many people outside of England have never heard of either of these, so here’s a bit of information on them.

Both Rebelstar and Laser Squad were written by the same people, the man behind both being Julian Gollop. Rebelstar Raiders was the first game, released on the Spectrum around 1983-84, before being re-written and re-released as Rebelstar later on. This really was a fantastic game, way ahead of its time, with brilliant graphics and sound (well, for a machine with a built-in speaker underneath, the sound was good)! It was the first game I know of that used the “action points” system.

Laser Squad followed a few years later, finally making it to the Amiga around 1989. This was still a brilliant game, and introduced the line-of-sight feature, the explosive weaponry and the concept of missions. However, for some reason it just didn’t have the charm of its predecessor, or so I felt anyway.

I understand that Julian is working on Laser Squad 2 at the minute. I hope this game not only fills the gap between LS 1 and 2, but stands on its own. I’ve tried to make it as original as possible, while still obviously being “strongly influenced” by the other games.

Anyway, I think AoW has a lot to offer of its own, apart from the influences of these games. Most obviously is the mission editor... it’s possible to do a lot more with this than you would think. Units don’t have to be people, they can be Battlecars, fighter planes, nuclear missiles... The three missions here are all sci-fi scenarios, but there’s no reason why they have to be. I know of someone who’s doing fantasy scenarios. Special Forces (SAS, Green Berets) are a possibility, and I’ve heard an unconfirmed rumour that someone is working on a Disney mission! I always wanted to see what the Chipmonks could do with a rocket launcher.

I have to mention Space Crusade by Gremlin here. That game has the honour of being the reason I wrote this. The game is OK, but I couldn’t help comparing it to Laser Squad. It contains several things I hate:

• Completely random death (“AutoDefences”... aargh!)

• You can stand right next to an enemy, shoot him with, like, a huge Autocannon, and miss him! How realistic. It wasn’t just that, it’s the way that the shot doesn’t hit anything else, it just disappears if it misses the assigned target. Ironically, you can still miss someone from right next to them in this game, but at least it’s a) unlikely, and b) fun when the shot hits his team member next to him!

• Randomly-appearing aliens. You get the feeling that, as the game realises it’s losing, it just rolls up another alien to throw at you.

These features are all the result of the game being a conversion of a boardgame, it’s inherently limited. It’s still a decent game, though I think the random deaths spoil it.

10 The End

Wow! You didn’t read all that rubbish, did you? You’re dedicated! Or very bored.

I always like doc files that rabbit on, like Jeff Minter’s epics. I suppose it sort of gives you an insight into the person that wrote the thing, rather than the “Press Z to go left” sort of file.

I’m going to stop here before I put the last of the audience to sleep!

Thanks for taking the trouble to look at the game, I hope you like it. I wrote it rather than get a job over the summer, so somebody better appreciate it!

Feel free to spread AoW around as much as you like - it’s in the Public Domain, though I keep the copyright. Upload it to bulletin boards, send it abroad, copy it for your friends.

Play the game, and get your chequebook out!

Dave Smith 2.49 pm September 1st 1992